Thad, a single taxpayer, has taxable income before the QBI deduction of $190,700. Thad, a CPA, operates an accounting practice as a single-member LLC (which he reports as a sole proprietorship). During 2019, his proprietorship generates a qualified business income of $150,000, W–2 wages of $125,000, and $10,000 of qualified property.

Assume the QBI amount is net of the self-employment tax deduction.

What is Thad's qualified business income deduction?
Answer #1

Answer:

Taxable income before the QBI deduction = $190,700.

This amount is above threshold amount of $160,700 but below $210,700

Applicable percentage = 100% - (190700 - 160700) / 50000 = 100% - 60% = 40%

Thad's qualified business income deduction = 150000 * 20% * 40% = $12,000

Thad's qualified business income deduction = $12,000